# SiNGER: A Clearer Voice DistillsVision Transformers Further

### v1

@@ Abstract
**footnotetext: These authors contributed equally to this work.
@@ 2 Related Works
Vision Transformers. ViTs (vit) underpin many VFMs and have become the representative architecture for large-scale visual learning. Unlike convolutional networks (resnet; mobilenet), ViTs rely on self-attention with fully connected layers. Since ViTs form the architectural core of most VFMs, studying them provides representative insights that generalize broadly. Similarly, parameter-efficient tuning methods such as LoRA (hu2021lora) highlight how minimal perturbations can effectively adapt VFMs, a perspective that motivates our artifact-suppressing perturbations. However, the quadratic complexity of the ViT architecture limits the practicality of large models despite their advanced representation.
@@
Artifacts in Transformers. Artifacts are a recurring issue in transformer models, degrading the representation quality. vit-needs-registers demonstrated that ViTs produce high-norm artifacts, particularly in background regions, harming interpretability and dense prediction. Practical suppression strategies include register tokens (vit-needs-registers), and recent work argues these artifacts arise from power-method-like accumulation across residual layers, aligning tokens with the leading left singular vector  (gandikota2023sinder). In the knowledge distillation domain, ViTKD (kd-vitkd) randomly masks teacher features to reduce the mimicking of high-norm artifacts. However, such indiscriminate masking also removes informative inlier signals, motivating artifact-aware KD that suppresses high-norm artifacts while preserving inlier structure.

### v3

@@ Abstract
[…]
@@ 2 Related Works
Vision Transformers. Visual transformers (vit; deit3) underpin many VFMs and have become the representative architecture for large-scale visual learning. Unlike convolutional networks (resnet; mobilenet), ViTs rely on self-attention with fully connected layers. Since ViTs form the architectural core of most VFMs, studying them provides representative insights that generalize broadly. Similarly, parameter-efficient tuning methods such as LoRA (hu2021lora) highlight how minimal perturbations can effectively adapt VFMs, a perspective that motivates our artifact-suppressing perturbations. However, the quadratic complexity of the ViT architecture limits the practicality of large models despite their advanced representation.
@@
Artifacts in Transformers. Artifacts are a recurring issue in visual transformer models, degrading the representation quality. vit-needs-registers demonstrated that ViTs produce high-norm artifacts, particularly in background regions, harming interpretability and dense prediction. Practical suppression strategies include register tokens (vit-needs-registers), and recent work argues these artifacts arise from power-method-like accumulation across residual layers, aligning tokens with the leading left singular vector  (gandikota2023sinder). In the knowledge distillation domain, ViTKD (kd-vitkd) randomly masks teacher features to reduce the mimicking of high-norm artifacts. However, such indiscriminate masking also removes informative inlier signals, motivating artifact-aware KD that suppresses high-norm artifacts while preserving inlier structure.